OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (5,6) and (9,14)?

OpenStudy (academicgurusinc):

First let us find the slope of the line using the two points (5,6), (9,14) given. x1 = 5, y1 = 6, x2 = 9, y2 =14 m = (y2 - y1)/(x2 - x1). m = (14 - 6)/(9 - 5). m = 8/4 = 2. Therefore slope, m = 2. Using y = mx + c. Slope = m, m = 2. y = 2x + c Substituting any of the point (5,6) into the equation, x = 5, y = 6. 6 = 2(5) + c 6 = 10 + c 6-10 = c -4 = c c = -4. Substituting into y = mx + c, m =2, c = -4 Therefore y = 2x + (-4). y = 2x - 4. I hope that helps. If you found this helpful, I encourage you to subscribe to our youtube channel, to stay current with all of our new videos: https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg You may also want to check out our Math Challenge contest, for a chance to win a $50 Visa gift card: https://www.youtube.com/watch?v=2Dbu-R_Hj7E Regards, Academic Gurus Inc.